OpenStudy (anonymous):

If A = {positive integers} and B = {...,−4, −3, −2, −1} what is A ∩ B?

OpenStudy (dan815):

A intersect B there will be no intersection

OpenStudy (dan815):

since positive integers are not part of the negative integers

OpenStudy (anonymous):

If you expand the entire equation, u should get the answer A^(6/2) * B^(6/3) = 432 A^3 * B^2 = 432 432 factors down to 16 * 27 16 = 4^2 and 27 = 3^3 So, A=3 and B=4 A*B = 3*4 AB = 12

OpenStudy (dan815):